For the cantilever beam and loading shown, use singularity functions or integration to determine the slope and deflection at6. PL2/24EI , PL3/48EI 1please use singularity functions
Answer #1

To I At get deflection, integrating eq 0 EI.Y = - w (1-x) - wsx + G the fixed end A x=0, y-o. L4 24 24 Hence, 24 putting x=